How many numbers can be made with the help of the digits0,1,2,3,4,5which are grater than 3000 (repetition is not all allowed

Grade:7

2 Answers

Sher Mohammad IIT Delhi
10 years ago
total number with 6 digit possible =6!=720
numbers less than or equal to 3000=3*5*4*3=180
total number which are greater than 3000= 720-180=540


sambit
10 years ago
no. of 4 digit no.s above 3000 is 3*5*4*3 =180 no. of 5 digit no.s = 5*5*4*3*2=600 no. of six digit no.s =5*5*4*3*2*1=600 total no. of required no.s = 180+600+600=1380 ..(q.e.d)